I am having a ton of trouble with this one please help...
1. Please show all of your work when answering the following:

I. Use Chebyshev’s theorem to find what percent of the values will fall between 201 and 345 for a data set with a mean of 273 and standard deviation of 18.

II. Use the Empirical Rule to find what two values 68% of the data will fall between for a data set with a mean of 216 and standard deviation of 15.

Answer by stanbon(75887)

. Use Chebyshev’s theorem to find what percent of the values will fall between 201 and 345 for a data set with a mean of 273 and standard deviation of 18.
---
z(201) = (201-273)/18 = -4
z(345) = (345-273)/18 = +4
-----
Percent of data is at least (1-(1/4)^2) = 1-(1/16) = 0.9375 = 93.75%
=====================================
II. Use the Empirical Rule to find what two values 68% of the data will fall between for a data set with a mean of 216 and standard deviation of 15.
----
Find the z-value with a left tail of 0.5-0.34 = 0.16
invNorm(0.16) = -0.9945
------
Find the Data Values:
x = zs+u
x = -0.9945*15 + 216 = 201.08
x = 0.9945*15 + 216 = 230.92
===================================
Cheers,
Stan H.
===============